# Transversely projective structures on Smooth Foliations on Surfaces

### v1

@@ Abstract.
Brunella’s classification implies that every smooth foliation on a compact complex surface admits a singular transversely projective structure. However, Biswas and Sorin’s recent work shows that certain foliations on compact complex surfaces, despite possessing a singular transversely projective structure, do not admit a regular transversely projective structure. Here, we describe the smooth foliations on compact complex surfaces which fail to possess a regular transversely projective structure.
@@ 1. Introduction
In [biswas2024holomorphic], Biswas and Sorin employed a dimension counting argument to establish that a sufficiently general smooth foliation on the product of elliptic curves does not admit a transversely projective structure. Inspired by their work, our pourpose here is to describe all smooth holomorphic foliations on compact complex surfaces which do not admit a transversely projective structure.
Naturally, we build on Brunella’s classification of smooth foliation on compact complex surfaces. According to [brunella-97-MR1474805], if $X$ is a compact Kähler surface and $\mathcal{F}$ is a smooth foliation on it then $\mathcal{F}$ is a fibration, or $\mathcal{F}$ is transverse to the fibers of a $\mathbb{P}^{1}$-bundle, i.e. $\mathcal{F}$ is a Riccati foliation, or $\mathcal{F}$ is transverse to the general fiber of an elliptic fibration, i.e. $\mathcal{F}$ is a turbulent foliation, or $\mathcal{F}$ is a linear foliation on a complex torus, or $\mathcal{F}$ is a transversely hyperbolic foliation with dense leaves. Moreover, if $X$ is a non-Kähler compact surface carrying a smooth foliation $\mathcal{F}$ then $X$ is a Hopf surface or an Inoue surface and $\mathcal{F}$ is any foliation on them.
@@
As already pointed out by Biswas and Sorin, turbulent foliations are not necessarily transversely projective. In Section 4 below, we characterize which turbulent foliations admit transverse projective structures, see for instance Theorem 4.4.
@@ Acknowledgements
We are indebt to Sorin Dumitrescu for bringing the problem addressed in this article to our attention and for correspondence regarding it. Furthermore, the authors express their appreciation for the financial support provided by CAPES/COFECUB and CNPq Projeto Universal 408687/2023-1 "Geometria das Equações Diferenciais Algébricas". Pereira acknowledges the support from CNPq (Grant number 301683/2019-0), and FAPERJ (Grant number E26/200.550/2023). Fazoli acknowledges the support of FAPERJ (Grant number E26/201.353/2023). Melo acknowledges the support of CAPES-PROEX.
@@ 2.2. Transversely projective structures
We recall one of the many equivalent definitions of transversely projective foliations, see for instance [LORAY_2007].
@@
    a holomorphic section $\sigma:X\rightarrow P$ transverse to $\mathcal{G}$
such that $\mathcal{F}=\sigma^{*}\mathcal{G}$. The datum $\mathcal{P}=(\pi,\mathcal{G},\sigma)$ is called a (smooth) transversely projective structure of $\mathcal{F}$.
@@
If $X$ is a curve and $\mathcal{F}$ is the foliation with $T_{\mathcal{F}}=0$ and $N_{\mathcal{F}}=T_{X}$ then a transversely projective structure for $\mathcal{F}$ is nothing but a projective structure for the Riemann surface $X$ as explained in [loray-martin-09-MR2647972, Section 1.5].
@@
We recall that the set of projective structures on a Riemann surface $X$ is parameterized by the affine space $H^{0}(X,(\Omega^{1}_{X})^{\otimes 2})$, see [loray-martin-09-MR2647972]. We recall below an analogue for transversely projective foliations of this classical fact.
@@
We refer to [biswas2024holomorphic] for a proof of Proposition 2.3.
@@ 2.3. Distinguished local first integrals
Definitions 2.1 and 2.4 are equivalent. To pass from the first to the second, one starts with a sufficiently fine open covering $\{U_{i}\}$ of $X$, consider $\mathbb{P}^{1}$-valued submersions $g_{i}:\pi^{-1}(U_{i})\to\mathbb{P}^{1}$ which are first integrals for ${\left.\kern-1.2pt\mathcal{G}\vphantom{\big{|}}\right|_{\pi^{-1}(U_{i})}}$ (they exist and any two differ by left composition by an element of $\operatorname{Aut}(\mathbb{P}^{1})$, because of the transversality between $\mathcal{G}$ and $\pi$), and consider the first integrals for ${\left.\kern-1.2pt\mathcal{F}\vphantom{\big{|}}\right|_{U_{i}}}$ defined by $g_{i}\circ\sigma$. Reciprocally, to pass from Definition 2.4 to Definition 2.1 one constructs the $\mathbb{P}^{1}$-bundle $P$ by gluing $U_{i}\times\mathbb{P}^{1}$ with $U_{j}\times\mathbb{P}^{1}$ using the transition function $f_{ij}$ and interpret the first integral $f_{i}$ as a section of the $\mathbb{P}^{1}$-bundle. For further details on the equivalence of the two definitions, the reader can consult [godbillon-91-MR1120547].
@@ 2.4. Canonical projective structures on Riemann surfaces
Following [loray-martin-09-MR2647972], we will call the projective structure given by Example 2.5, the canonical projective structure of $X$. Example 2.5 has two natural counterparts for codimension one foliations.
@@
Let $\mathcal{F}$ be a foliation defined by a smooth fibration $f:X\to B$, that is, every fiber of $f$ is smooth and none of them is multiple. Then $\mathcal{F}$ admits a smooth transversely projective structure.
@@
The transversality between $\mathcal{F}$ and $\pi$ impose strong restrictions on the foliation $\mathcal{F}$ and on the fibration $\pi$ as explained in [zbMATH03907374, Chapter 5] which we summarize. Lifting paths on the basis $B$ along leaves of $\mathcal{F}$ one verifies that any two fibers of $\pi$ are isomorphic. Let us choose one of them, say $F_{b}=\pi^{-1}(b)$. Moreover, if $\tilde{B}$ is the universal covering of $B$ and $\tilde{X}$ is the fiber product $X\times_{B}\tilde{B}$ then $\tilde{X}$ is biholomorphic to the product $\tilde{B}\times F_{b}$ in such a way that $\tilde{\mathcal{F}}$, the pull-back of $\mathcal{F}$ to $\tilde{X}$, is the foliation defined the projection on the second the factor. The original manifold $X$ can be recovered as a quotient of $\tilde{X}$ by $\pi_{1}(B,b)$ acting on $\tilde{B}\times F$ through deck transformations on the first factor and through a representation $\rho:\pi_{1}(B,b)\to\operatorname{Aut}(F_{b})$.
@@ 2.6. Singular projective structures
    a codimension one smooth foliation $\mathcal{G}$ on $P$ transverse to the generic fiber of $\pi$; and
@@
It is well-known, and easy to check using Brunella’s classification, that every smooth foliation on a compact complex surface admits a singular transversely projective structure. For more about singular transversely projective foliations/structures, we refer to [Scardua1997] (the first work to consider the concept), [LORAY_2007] for a detailed discussion of the above definition, and [loray2016representations] for a description of the structure of codimension one singular transversely projective foliations on projective manifolds of arbitrary dimension.
@@ 3.1. Orbifold base
and consider the pair $(B,\Delta)$ as the orbifold base of the fibration in the sense of [zbMATH02123576, Section 1].
@@
A standard computation, which traces back to Darboux, shows that $N^{*}_{\mathcal{F}}=f^{*}\omega_{B}\otimes\mathcal{O}_{X}(f^{*}\Delta)$.
An orbifold base $(B,\Delta)$ is called good if there exists a curve $C$ and a discrete subgroup $G\subset\operatorname{Aut}(C)$ such that $B\simeq C/G$ and the ramification divisor of the quotient morphism $\pi:C\to C/G\simeq B$ equals $\pi^{*}\Delta$. We will call the triple $(C,G,\pi)$ an unfolding of $(B,\Delta)$. Otherwise, the orbifold base $(B,\Delta)$ is called bad.
Most orbifolds are good as established by Bundgaard-Nielsen-Fox theorem stated below. For a brief account of its history, the reader can consult [uludag-07-MR2306159, Theorem 1.4]. We also refer to [zbMATH04046721] and [zbMATH00558225, Chapter 2] for more information about orbifolds.
@@ 3.3. Quasi-smooth fibrations on compact Kähler surfaces
Let $g:L\rightarrow L$ be a non-trivial deck transformation. By Lefschetz trace formula (see for instance [hatcher-02-MR1867354, Theorem 2C.3]),
$\sum_{i=0}^{2\dim L}(-1)^{i}\mathrm{tr}(g_{*}:H_{i}(L,\mathbb{C})\rightarrow H_{i}(L,\mathbb{C}))=0.$
@@
$\sum_{i=0}^{2\dim L}(-1)^{i}\mathrm{tr}(g_{*}:H_{i}(L,\mathbb{C})\rightarrow H_{i}(L,\mathbb{C}))=\sum_{i=0}^{2\dim L}(-1)^{i}h_{i}(L,\mathbb{C})=\chi_{top}(L)$
@@
Let $\iota:L\rightarrow X$ be the inclusion of the general fiber $L$ of $f$. According to Deligne’s invariant subspace theorem, see for instance [zbMATH01927232, Theorem 16.2], the image of $\iota^{*}:H^{1}(X,\mathbb{C})\rightarrow H^{1}(L,\mathbb{C})$ is the subspace invariant by the monodromy representation of the local system $R^{1}f_{*}\mathbb{C}$ defined on $B^{\circ}$. Because it is trivial, we conclude that $\iota^{*}$ is surjective. Furthermore, $\iota^{*}$ defines a morphism of Hodge structures. Hence we get a surjective morphism
@@
Let $\omega\in H^{0}(X,\Omega^{1}_{X})$ such that $\iota^{*}\omega=\omega_{L}$ and consider the cohomology class of $\alpha=[\omega\wedge\overline{\omega}]\in H^{2}(X,\mathbb{C})$. Note that the contraction of $\alpha$ with the homology class of $L$ is equal to $\int_{L}\alpha$ and, therefore, non-zero. It follows that the same holds true for every other fiber of $f$ (multiple or not) since their homology classes all lie on the same line. Hence the $1$-form $\omega$ is non-zero in every fiber and it defines a foliation $\mathcal{G}$ completely transversal to $\mathcal{F}$. In particular, $T_{X}=T_{\mathcal{F}}\oplus T_{\mathcal{G}}$. Using [beauville-MR1760872, Theorem C], see also [zbMATH05145488] and [zbMATH07634625, Theorem 1.4], the universal cover of $\pi:\widetilde{X}\rightarrow X$ is the product of two simply-connected Riemann surfaces, and the foliation $\pi^{*}(\mathcal{F})$ is the fibration over one of them. This would imply that $B$ is a good orbifold, a contradiction. ∎
@@ 3.5. Quasi-smooth fibrations on projective threefolds
After applying the minimal model program (MMP) for codimension one smooth foliations on projective threefolds [zbMATH07132579, Theorem 1.3] we can assume that every fiber of $f$ is minimal, i.e. does not contain a smooth curve isomorphic to $\mathbb{P}^{1}$ with self-intersection $-1$.
@@
Lemma 3.6 implies that every fiber of $f$ has zero topological Euler characteristic. An inspection of the Enriques-Kodaira classification [bpv-MR2030225, Chapter VI, Table 10] shows that the fibers of $f$ are
@@
If the fibers are abelian surfaces then we can argue as in the proof of Theorem 3.4, but now using $2$-forms instead of $1$-forms, to obtain a one-dimensional foliation $\mathcal{G}$ everywhere transverse to $\mathcal{F}$. It follows from [zbMATH05145488] that the universal covering is the product $\mathbb{C}^{2}\times B=\mathbb{C}^{2}\times\mathbb{P}^{1}$ with $f$ lifting to the fibration defined by projection to $\mathbb{P}^{1}$. It follows that $f$ is good.
If the fibers are bi-elliptic surfaces or have Kodaira dimension one then there exists a smooth foliation $\mathcal{E}$ by elliptic curves tangent to the fibration. Since $\mathcal{E}$ is smooth, it must be an isotrivial fibration. Therefore we can apply [zbMATH07634625, Theorem 1.4] to produce a finite étale covering $p:Y\to X$ such that $p^{*}\mathcal{E}$ is a locally trivial elliptic fibration (no multiple fibers). Again, the fibration $f$ descends to a fibration on a smooth projective surface and we can conclude using 3.4 as in the case of negative Kodaira dimension . ∎

### v2

@@ Abstract.
Brunella’s classification implies that every smooth foliation on a compact complex surface admits a singular transversely projective structure. However, Biswas and Dumitrescu’s recent work shows that certain foliations on compact complex surfaces, despite possessing a singular transversely projective structure, do not admit a regular transversely projective structure. Here, we describe the smooth foliations on compact complex surfaces which fail to possess a regular transversely projective structure.
@@ 1. Introduction
In [3], Biswas and Dumitrescu employed a dimension counting argument to establish that a sufficiently general smooth foliation on the product of elliptic curves does not admit a transversely projective structure. Inspired by their work, our pourpose here is to describe all smooth holomorphic foliations on compact complex surfaces which do not admit a transversely projective structure.
Naturally, we build on Brunella’s classification of smooth foliations on compact complex surfaces. According to [4], if $X$ is a compact Kähler surface and $\mathcal{F}$ is a smooth foliation on it then $\mathcal{F}$ is a fibration, or $\mathcal{F}$ is transverse to the fibers of a $\mathbb{P}^{1}$-bundle, i.e. $\mathcal{F}$ is a Riccati foliation, or $\mathcal{F}$ is transverse to the general fiber of an elliptic fibration, i.e. $\mathcal{F}$ is a turbulent foliation, or $\mathcal{F}$ is a linear foliation on a complex torus, or $\mathcal{F}$ is a transversely hyperbolic foliation with dense leaves. Moreover, if $X$ is a non-Kähler compact surface carrying a smooth foliation $\mathcal{F}$ then $X$ is a Hopf surface or an Inoue surface and $\mathcal{F}$ is any foliation on them.
@@
As already pointed out by Biswas and Dumitrescu, turbulent foliations are not necessarily transversely projective. In Section 4 below, we characterize which turbulent foliations admit transverse projective structures, see for instance Theorem 4.4.
@@ Acknowledgements
We are indebt to Sorin Dumitrescu for bringing the problem addressed in this article to our attention and for correspondence regarding it. Furthermore, the authors express their appreciation for the financial support provided by: CAPES-COFECUB program (project number: Ma 1017/24, Géométrie des équations différentielles et des variétés algébriques), funded by the French Ministry for Europe and Foreign Affairs, the French Ministry for Higher Education and Research, and CAPES; and CNPq Projeto Universal 408687/2023-1 "Geometria das Equações Diferenciais Algébricas". Pereira acknowledges the support from CNPq (Grant number 301683/2019-0), and FAPERJ (Grant number E26/200.550/2023). Fazoli acknowledges the support of FAPERJ (Grant number E26/201.353/2023). Melo acknowledges the support of CAPES-PROEX.
@@ 2.2. Transversely projective structures
We recall one of the many equivalent definitions of transversely projective foliations, see for instance [16].
@@
    a holomorphic section $\sigma:X\rightarrow P$ transverse to $\mathcal{G}$ at every point, such that $\mathcal{F}=\sigma^{*}\mathcal{G}$.
The datum $\mathcal{P}=(\pi,\mathcal{G},\sigma)$ is called a (smooth) transversely projective structure of $\mathcal{F}$.
@@
If $X$ is a curve and $\mathcal{F}$ is the foliation with $T_{\mathcal{F}}=0$ and $N_{\mathcal{F}}=T_{X}$ then a transversely projective structure for $\mathcal{F}$ is nothing but a projective structure for the Riemann surface $X$ as explained in [15, Section 1.5].
@@
The set of projective structures on a Riemann surface $X$ is parameterized by the affine space $H^{0}(X,(\Omega^{1}_{X})^{\otimes 2})$, see [15]. We recall below an analogue for transversely projective foliations of this classical fact.
@@
We refer to [3] for a proof of Proposition 2.3.
@@ 2.3. Distinguished local first integrals
Definitions 2.1 and 2.4 are equivalent. To pass from the first to the second, one starts with a sufficiently fine open covering $\{U_{i}\}$ of $X$, consider $\mathbb{P}^{1}$-valued submersions $g_{i}:\pi^{-1}(U_{i})\to\mathbb{P}^{1}$ which are first integrals for ${\left.\kern-1.2pt\mathcal{G}\vphantom{\big{|}}\right|_{\pi^{-1}(U_{i})}}$ (they exist and any two differ by left composition by an element of $\operatorname{Aut}(\mathbb{P}^{1})$, because of the transversality between $\mathcal{G}$ and $\pi$), and consider the first integrals for ${\left.\kern-1.2pt\mathcal{F}\vphantom{\big{|}}\right|_{U_{i}}}$ defined by $g_{i}\circ\sigma$. Reciprocally, to pass from Definition 2.4 to Definition 2.1 one constructs the $\mathbb{P}^{1}$-bundle $P$ by gluing $U_{i}\times\mathbb{P}^{1}$ with $U_{j}\times\mathbb{P}^{1}$ using the transition function $f_{ij}$ and interpret the first integral $f_{i}$ as a section of the $\mathbb{P}^{1}$-bundle. For further details on the equivalence of the two definitions, the reader can consult [12].
@@ 2.4. Canonical projective structures on Riemann surfaces
Following [15], we will call the projective structure given by Example 2.5, the canonical projective structure of $X$. Example 2.5 has two natural counterparts for codimension one foliations.
@@
Let $B$ be a Riemann surface. Let $\mathcal{F}$ be a foliation defined by a smooth fibration $f:X\to B$, that is, every fiber of $f$ is smooth and none of them is multiple. Then, every projective structure for $B$ induces a smooth transversely projective structure for $\mathcal{F}$, given as the pullback of the projective atlas. Since, by Example 2.5, every Riemann surface admits a smooth projective structure, then $\mathcal{F}$ is transversely projective.
@@
The transversality between $\mathcal{F}$ and $\pi$ impose strong restrictions on the foliation $\mathcal{F}$ and on the fibration $\pi$ as explained in [7, Chapter 5] which we summarize. Lifting paths on the basis $B$ along leaves of $\mathcal{F}$ one verifies that any two fibers of $\pi$ are isomorphic. Let us choose one of them, say $F_{b}=\pi^{-1}(b)$. Moreover, if $\tilde{B}$ is the universal covering of $B$ and $\tilde{X}$ is the fiber product $X\times_{B}\tilde{B}$ then $\tilde{X}$ is biholomorphic to the product $\tilde{B}\times F_{b}$ in such a way that $\tilde{\mathcal{F}}$, the pull-back of $\mathcal{F}$ to $\tilde{X}$, is the foliation defined the projection on the second the factor. The original manifold $X$ can be recovered as a quotient of $\tilde{X}$ by $\pi_{1}(B,b)$ acting on $\tilde{B}\times F$ through deck transformations on the first factor and through a representation $\rho:\pi_{1}(B,b)\to\operatorname{Aut}(F_{b})$.
@@ 2.6. Singular projective structures
    a codimension one foliation $\mathcal{G}$ on $P$ transverse to the generic fiber of $\pi$; and
@@
It is well-known, and easy to check using Brunella’s classification, that every smooth foliation on a compact complex surface admits a singular transversely projective structure. For more about singular transversely projective foliations/structures, we refer to [18] (the first work to consider the concept), [16] for a detailed discussion of the above definition, and [17] for a description of the structure of codimension one singular transversely projective foliations on projective manifolds of arbitrary dimension.
@@ 3.1. Orbifold base
and consider the pair $(B,\Delta)$ as the orbifold base of the fibration in the sense of [8, Section 1].
@@
A standard computation, which traces back to Darboux, shows that $N^{*}_{\mathcal{F}}=f^{*}\omega_{B}\otimes\mathcal{O}_{X}(f^{*}\Delta)$ (see [5, Section 2.3, Example 5]).
An orbifold base $(B,\Delta)$ is called good (also called uniformizable) if there exists a curve $C$ and a discrete subgroup $G\subset\operatorname{Aut}(C)$ such that $B\simeq C/G$ and the ramification divisor of the quotient morphism $\pi:C\to C/G\simeq B$ equals $\pi^{*}\Delta$. We will call the triple $(C,G,\pi)$ an unfolding (also called uniformization) of $(B,\Delta)$. Otherwise, the orbifold base $(B,\Delta)$ is called bad.
Most orbifolds are good as established by Bundgaard-Nielsen-Fox theorem stated below. For a brief account of its history, the reader can consult [20, Theorem 1.4]. We also refer to [14] and [10, Chapter 2] for more information about orbifolds.
@@ 3.3. Quasi-smooth fibrations on compact Kähler surfaces
Let $g:L\rightarrow L$ be a non-trivial deck transformation. By Lefschetz trace formula (see for instance [13, Theorem 2C.3]),
$\sum_{i=0}^{2\dim L}(-1)^{i}\mathrm{tr}\Big{(}g_{*}:H_{i}(L,\mathbb{C})\rightarrow H_{i}(L,\mathbb{C})\Big{)}=0.$
@@
$\sum_{i=0}^{2\dim L}(-1)^{i}\mathrm{tr}\Big{(}g_{*}:H_{i}(L,\mathbb{C})\rightarrow H_{i}(L,\mathbb{C})\Big{)}=\sum_{i=0}^{2\dim L}(-1)^{i}h_{i}(L,\mathbb{C})=\chi_{top}(L)$
@@
Let $\iota:L\rightarrow X$ be the inclusion of the general fiber $L$ of $f$. According to Deligne’s invariant subspace theorem, see for instance [21, Theorem 16.2], the image of $\iota^{*}:H^{1}(X,\mathbb{C})\rightarrow H^{1}(L,\mathbb{C})$ is the subspace invariant by the monodromy representation of the local system $R^{1}f_{*}\mathbb{C}$ defined on $B^{\circ}$. Because it is trivial (see proof of Lemma 3.6), we conclude that $\iota^{*}$ is surjective. Furthermore, $\iota^{*}$ defines a morphism of Hodge structures. Hence we get a surjective morphism
@@
Let $\omega\in H^{0}(X,\Omega^{1}_{X})$ such that $\iota^{*}\omega=\omega_{L}$ and consider the cohomology class of $\alpha=[\omega\wedge\overline{\omega}]\in H^{2}(X,\mathbb{C})$. Note that the contraction of $\alpha$ with the homology class of $L$ is equal to $\int_{L}\alpha$ and, therefore, non-zero. It follows that the same holds true for every other fiber of $f$ (multiple or not) since their homology classes all lie on the same line. Hence the $1$-form $\omega$ is non-zero in every fiber and it defines a foliation $\mathcal{G}$ completely transversal to $\mathcal{F}$. In particular, $T_{X}=T_{\mathcal{F}}\oplus T_{\mathcal{G}}$. Using [2, Theorem C], see also [6] and [9, Theorem 1.4], the universal cover of $\pi:\widetilde{X}\rightarrow X$ is the product of two simply-connected Riemann surfaces, and the foliation $\pi^{*}(\mathcal{F})$ is the fibration over one of them. This would imply that $B$ is a good orbifold, a contradiction. ∎
3.5. Quasi-smooth fibrations on surfaces with bad orbifold base
From the Theorem 3.4, we know that if $f:X\rightarrow\mathbb{P}^{1}$ is a quasi-smooth fibration with bad orbifold base, then $X$ must be a non Kähler compact surface. Moreover, in Example 3.3, we constructed examples of quasi-smooth fibrations on Hopf surfaces with bad orbifold base. From this, one can ask if there are other examples. The next proposition shows that this is the only case.
Proposition 3.7.
Let $X$ be a compact complex surface and $f:X\rightarrow\mathbb{P}^{1}$ be a quasi-smooth fibration with bad orbifold base. Then $X$ is a Hopf surface.
Proof.
Let $f:X\rightarrow\mathbb{P}^{1}$ be a fibration with a bad orbifold base $(\mathbb{P}^{1},\frac{m-1}{m}p+\frac{n-1}{n}q)$ (for the case $(\mathbb{P}^{1},\frac{m-1}{m}p)$ the computation is similar). By Lemma 3.6, $f$ is an elliptic fibration, and since the fibers are irreducible curves, they can not be $(-1)$-curves. That is, $f$ is a relatively minimal elliptic fibration.
The first step is to conclude that $\operatorname{kod}(X)=-\infty$. By [1, Chapter V, Proposition 12.5], $\operatorname{kod}(X)=-\infty$ if, and only if, $\delta(f)<0$, where $\delta(f)$ of a elliptic fibration $f:X\rightarrow B$ is defined as
$\delta(f):=\chi(X,\mathcal{O}_{X})+\left(2g(S)-2+\deg(\Delta)\right).$
In our situation, we have that $\chi(X,\mathcal{O}_{X})=deg(f_{*}\omega_{X|\mathbb{P}^{1}})$ by [1, Chapter III, Theorem 18.2], and by [1, Chapter III, Theorem 18.2], we have $deg(f_{*}\omega_{X|\mathbb{P}^{1}})=0$. Therefore,
$\delta(f)=\chi(X,\mathcal{O}_{X})-\frac{1}{m}-\frac{1}{n}=-\frac{1}{m}-\frac{1}{n}<0,$
and we conclude that $\operatorname{kod}(X)=-\infty$.
From Theorem 3.4 and Lemma 3.6, we see that $X$ must be a non Kähler surface. By [1, Chapter V, Theorem 18.6], it follows that $X$ is a Hopf surface.
@@ 3.6. Quasi-smooth fibrations on projective threefolds
After applying the minimal model program (MMP) for codimension one smooth foliations on projective threefolds [19, Theorem 1.3] we can assume that every fiber of $f$ is minimal, i.e. does not contain a smooth curve isomorphic to $\mathbb{P}^{1}$ with self-intersection $-1$.
@@
Lemma 3.6 implies that every fiber of $f$ has zero topological Euler characteristic. An inspection of the Enriques-Kodaira classification [1, Chapter VI, Table 10] shows that the fibers of $f$ are
@@
If the fibers are abelian surfaces then we can argue as in the proof of Theorem 3.4, but now using $2$-forms instead of $1$-forms, to obtain a one-dimensional foliation $\mathcal{G}$ everywhere transverse to $\mathcal{F}$. It follows from [6] that the universal covering is the product $\mathbb{C}^{2}\times B=\mathbb{C}^{2}\times\mathbb{P}^{1}$ with $f$ lifting to the fibration defined by projection to $\mathbb{P}^{1}$. It follows that $f$ is good.
If the fibers are bi-elliptic surfaces or have Kodaira dimension one then there exists a smooth foliation $\mathcal{E}$ by elliptic curves tangent to the fibration. Since $\mathcal{E}$ is smooth, it must be an isotrivial fibration. Therefore we can apply [9, Theorem 1.4] to produce a finite étale covering $p:Y\to X$ such that $p^{*}\mathcal{E}$ is a locally trivial elliptic fibration (no multiple fibers). Again, the fibration $f$ descends to a fibration on a smooth projective surface and we can conclude using 3.4 as in the case of negative Kodaira dimension . ∎